# Spike No More: Stabilizing the Pre-training of Large Language Models

### v3

@@ 7 Related Work
To stabilize trainings of Transformer-based neural language models, there have been various discussions on the architecture (Xiong et al., 2020; Liu et al., 2020; Takase et al., 2023; Zeng et al., 2023; Zhai et al., 2023), initialization method (Nguyen & Salazar, 2019; Zhang et al., 2019; Huang et al., 2020; Wang et al., 2022), training strategy (Zhang et al., 2022; Li et al., 2022), and loss function (Chowdhery et al., 2022; Wortsman et al., 2023). Xiong et al. (2020) theoretically analyzed gradient scales of each part in Transformers, and indicated that the Pre-LN Transformer is more stable than the Post-LN Transformer, that is the original Transformer architecture (Vaswani et al., 2017). Since the Pre-LN Transformer is more stable than the Post-LN Transformer theoretically and empirically, recent studies mainly have used the Pre-LN Transformer to construct an LLM. We also assume using the Pre-LN Transformer in the analysis on the training dynamics in this paper. To stabilize the LLM pre-training, Le Scao et al. (2022) applied the layer normalization to the embedding layer. Zeng et al. (2023) used shrink embedding gradient technique (Ding et al., 2021). In this study, we theoretically proved that the layer normalization to the embedding layer controls the upper bound of the gradient norms of sub-layers when we use the widely used initialization method for LLMs (Nguyen & Salazar, 2019; Shoeybi et al., 2020), and thus, it stabilizes the pre-training. For the initialization methods, Nguyen & Salazar (2019) proposed a strategy to initialize parameters of Transformers with small values to stabilize their training. Zhang et al. (2019) and Huang et al. (2020) indicated that we can remove layer normalizations in Transformers if we use their proposed initialization methods. Wang et al. (2022) adjusted initial parameter scales based on the number of layers to stabilize the Post-LN Transformer. In this study, we indicated that the widely used initialization method (Shoeybi et al., 2020), which makes parameters small, is necessary to stabilize the LLM pre-training. Moreover, we proved that we can prevent the loss spike problem by making the standard deviation of embeddings close to $1$.
@@
As shown in Table 3, our modification enables the pre-training with a relatively larger learning rate, and can achieve better performance. Thus, this study can be regarded as on the efficiency of LLM pre-training because our modification can construct a better LLM with a given budget. Strubell et al. (2019) and Schwartz et al. (2019) reported that recent neural methods require substantial computational costs, and thus, they argued that we have to explore a cost-efficient approach. Rajbhandari et al. (2020) proposed ZeRO that reduces memory redundancies during the multi GPU training without increasing communication volume. Dao et al. (2022) focused on GPU memory reads/writes, and proposed FlashAttention that accelerates the speed of attention mechanisms in Transformers. To reduce the number of computations in the attention mechanism, Shazeer (2019) proposed the multi-query attention that shares one key and value across all of the attention heads in each layer. Takase & Kiyono (2023) explored several parameter sharing strategies, and indicated that parameter sharing across some layers can achieve comparable performance to the vanilla model with a small number of parameters. Moreover, several studies have explored a better construction way with a limited budget (Izsak et al., 2021; Takase & Kiyono, 2021). We believe that we can take advantage of their findings to make our LLMs more efficient.
@@ 8 Conclusion
This paper explored why large language models (LLMs) sometimes experience loss spikes during pre-training. To provide evidence, we specifically focused on the gradients of sub-layers. We introduced an upper bound for the gradient norms through an analysis of the spectral norms of the Jacobian matrices for the sub-layers. We then theoretically identified two conditions for avoiding loss spikes: small sub-layers and large shortcut. To meet these conditions, we show that using the widely adopted initialization method for LLMs can make the sub-layer parameters small, and that embedding scaling or incorporating layer normalization into the embedding layer can make the standard deviation of each embedding close to $1$, resulting in large shortcut. Experimental results indicated that methods satisfying these conditions avoid loss spikes. Furthermore, these methods allow for training with a relatively larger learning rate, leading to improved performance. We hope our theoretical analyses and empirical findings will help avoid wasting valuable time and computational budgets during LLM construction.
@@
To stabilize the LLM pre-training, this paper provides theoretical analyses om the spectral norms of the Jacobian matrices for sub-layers to estimate the upper bound of the gradient norm of $\mathcal{L}$. This paper focuses on only the stability of LLM pre-training, and thus, we have to address other issues of LLMs such as hallucinations to use the LLM in a real application.
@@
We do not aim to propose a novel method in this paper, but we mainly focus on theoretical analyses on the spectral norms of the Jacobian matrices to find the factor to stabilize the pre-training of LLMs. We justify our theoretical analyses through experiments with various situations. To activate our modification, we add only several lines to a widely used implementation, i.e., Megatron-LM [^12] https://github.com/NVIDIA/Megatron-LM. Therefore, we believe that it is easy to reproduce our experimental results. However, because it is difficult to conclude that our provided conditions completely solve the instability during pre-training of LLMs, it is better to combine other techniques to stabilize the pre-training such as an auxiliary loss described by Chowdhery et al. (2022) to make the pre-training more stable in an actual pre-training situation.
@@ Appendix A Hyper-Parameters
Table 4 shows that hyper-parameters used in our experiments on LLMs. In addition to experiments described in Section 5, this table also indicates the hyper-parameters of the model with 13B parameters that we evaluated in Appendix C.
@@
Appendix B Methods without Small Sub-Layers
(a) 350M parameters.
(b) 1.7B parameters.
(a) 350M parameters.
(b) 1.7B parameters.
Table 5: Perplexities of each method.
@@ Appendix B Methods without Small Sub-Layers
Appendix C Pre-training of the Model with 13B Parameters
(a) $\mathrm{lr}=3.0\times 10^{-4}$
(b) $\mathrm{lr}=1.0\times 10^{-4}$.
@@ Appendix C Pre-training of the Model with 13B Parameters
Vanilla N/A 15.12 N/A 6.50
Scaled Embed 14.47 15.25 5.97 6.53
We conducted pre-trainings of models with 13B parameters to indicate that our modification can stabilize a model with many more parameters than the ones discussed in Section 5. To make this experiment close to a realistic situation, as shown in Table 4, we increased the batch size and the number of updates, and decreased the Adam $\beta_{2}$. In particular, for Adam $\beta_{2}$, most studies have used $0.95$ to stabilize their pre-trainings (Brown et al., 2020; Zhang et al., 2022; Zeng et al., 2023; Biderman et al., 2023; Touvron et al., 2023), and thus, we also used $0.95$ in this experiment. We tried two learning rates: $3.0\times 10^{-4}$, which is the same value in Touvron et al. (2023), and $1.0\times 10^{-4}$.
Figure 9 shows the loss values of each configuration in validation data. As shown in (a) of this figure, the loss value of Vanilla rose from approximately 10000 steps in $\mathrm{lr}=3.0\times 10^{-4}$. Then, the gradient of this model became too large to continue its pre-training. In contrast, the loss value of Scaled Embed consistently decreased. This result indicates that Scaled Embed stabilized the pre-training. We emphasize that the pre-training of Vanilla is essentially unstable even if we use the widely used Adam $\beta_{2}$ value, $0.95$, which is known as the technique to stabilize the pre-training, and our modification is also effective for the stabilization in this realistic situation.
Table 6 shows the perplexities of each configuration in evaluation data. This table indicates that we can achieve better performance when we use a larger learning rate in the same as in Section 6.1. In addition, the perplexities of Scaled Embed were comparable to ones of Vanilla when we used the small learning rate: $\mathrm{lr}=1.0\times 10^{-4}$. These results imply that our modification has no considerable risk in pre-training. Thus, we have to satisfy large shortcut in addition to small sub-layers to stabilize the pre-trainings of LLMs.
Appendix D RMSNorm
@@
Figure 10 shows the loss values and gradient norms of the Vanilla configuration in Section 5 and the one using RMSNorms instead of LNs (“RMSNorm” in figures) with 350M parameters. We trained them with $\mathrm{lr}=5.0\times 10^{-4}$ as in Section 5131313We tried to train them with $\mathrm{lr}=1.0\times 10^{-3}$ but RMSNorm exploded.. As shown in these figures, RMSNorm faced loss and gradient spikes in a similar location to the ones of Vanilla. These empirical results also indicate that the RMSNorms have the same problem as LNs regarding the instability.
@@
Appendix E Scaling Embeddings with Larger Value
Equations (15) and (20) indicate that we can stabilize the LLM pre-training by adjusting the standard deviation of the shortcut to a large value. In fact, our experimental results show that we can stabilize the LLM pre-training by making the standard deviation of each embedding close to $1$. To investigate how about a larger value, we conducted experiments with making the standard deviation of each embedding close to 5 and 50.
Figure 11 shows loss curves in validation data for 350M and 1.7B parameters in each situation. This figure indicates that although all settings prevented the loss spike problem, the larger standard deviation than $1$ degraded the performance. Therefore, it is unnecessary to scale the standard deviation of each embedding with a larger value than $1$ to prevent the performance degradation.
@@ Appendix E Scaling Embeddings with Larger Value
Figure 11: Loss curves in validation data when we scale the standard deviation of embeddings with larger than $1$.
Appendix F Distributions of Sub-Layer Outputs
Figure 12 shows output distributions of each sub-layer for each layer at the initialization. This figure indicates that each sub-layer output is close to the normal distribution in various configurations. Therefore, the assumption in this study, which is that the vector $x$ at each layer follows the normal distribution, is reasonable.
@@ Appendix F Distributions of Sub-Layer Outputs
Figure 12: Output distributions of each sub-layer.
Appendix G Discussion on ReLU in FFN

### v4

@@ 6 Related Work
To stabilize trainings of Transformer-based neural language models, there have been various discussions on the architecture (Xiong et al., 2020; Liu et al., 2020; Takase et al., 2023; Zeng et al., 2023; Zhai et al., 2023), initialization method (Nguyen & Salazar, 2019; Zhang et al., 2019; Huang et al., 2020; Wang et al., 2022), training strategy (Zhang et al., 2022; Li et al., 2022), and loss function (Chowdhery et al., 2022; Wortsman et al., 2023).
Xiong et al. (2020) theoretically analyzed gradient scales of each part in Transformers, and indicated that the Pre-LN Transformer is more stable than the Post-LN Transformer, that is the original Transformer architecture (Vaswani et al., 2017). Since the Pre-LN Transformer is more stable than the Post-LN Transformer theoretically and empirically, recent studies mainly have used the Pre-LN Transformer to construct an LLM. We also assume using the Pre-LN Transformer in the analysis on the training dynamics in this paper.
To stabilize the LLM pre-training, Le Scao et al. (2022) applied the layer normalization to the embedding layer. Zeng et al. (2023) used shrink embedding gradient technique (Ding et al., 2021). Nishida et al. (2024) proposed weight scaling as reparameterization (WeSaR) which uses additional parameters to scale parameters of internal layers. In this study, we theoretically proved that the layer normalization to the embedding layer controls the upper bound of the gradient norms of sub-layers when we use the widely used initialization method for LLMs (Nguyen & Salazar, 2019; Shoeybi et al., 2020), and thus, it stabilizes the pre-training.
For the initialization methods, recent studies have proposed maximal update parameterization ($\mu$P) and its variants as approaches to transfer hyper-parameters from small models to larger ones without incurring the cost of hyper-parameter search Yang & Hu (2021); Yang et al. (2024a; b). However, as described in Yang et al. (2024b), such methods may be ineffective when applied to Transformers, as their underlying assumptions do not accurately reflect the actual Transformer architectures. In studies focusing on Transformers, Nguyen & Salazar (2019) proposed a strategy to initialize parameters of Transformers with small values to stabilize their training. Zhang et al. (2019) and Huang et al. (2020) indicated that we can remove layer normalizations in Transformers if we use their proposed initialization methods. Wang et al. (2022) adjusted initial parameter scales based on the number of layers to stabilize the Post-LN Transformer. In this study, we indicated that the widely used initialization method (Shoeybi et al., 2020), which makes parameters small, is necessary to stabilize the LLM pre-training. Moreover, we proved that we can prevent the loss spike problem by making the standard deviation of embeddings close to 111.
@@
As shown in Section 5.4 and Appendix B.2, our modification enables the pre-training with a relatively larger learning rate, and can achieve better performance. Thus, this study can be regarded as on the efficiency of LLM pre-training because our modification can construct a better LLM with a given budget. Strubell et al. (2019) and Schwartz et al. (2019) reported that recent neural methods require substantial computational costs, and thus, they argued that we have to explore a cost-efficient approach. Rajbhandari et al. (2020) proposed ZeRO that reduces memory redundancies during the multi GPU training without increasing communication volume. Dao et al. (2022) focused on GPU memory reads/writes, and proposed FlashAttention that accelerates the speed of attention mechanisms in Transformers. To reduce the number of computations in the attention mechanism, Shazeer (2019) proposed the multi-query attention that shares one key and value across all of the attention heads in each layer. Takase & Kiyono (2023) explored several parameter sharing strategies, and indicated that parameter sharing across some layers can achieve comparable performance to the vanilla model with a small number of parameters. Moreover, several studies have explored a better construction way with a limited budget (Izsak et al., 2021; Takase & Kiyono, 2021). We believe that we can take advantage of their findings to make our LLMs more efficient.
@@ 7 Conclusion
This paper explored why large language models (LLMs) sometimes experience loss spikes during pre-training. To provide evidence, we specifically focused on the gradients of sub-layers. We introduced an upper bound for the gradient norms through an analysis of the spectral norms of the Jacobian matrices for the sub-layers. We then theoretically identified two conditions for avoiding loss spikes: small sub-layers and large shortcut. To meet these conditions, we show that using the widely adopted initialization method for LLMs can make the sub-layer parameters small, and that embedding scaling or incorporating layer normalization into the embedding layer can make the standard deviation of each embedding close to 111, resulting in large shortcut. Experimental results indicated that methods satisfying these conditions avoid loss spikes. Furthermore, these methods allow for training with a relatively larger learning rate, leading to improved performance. We hope our theoretical analyses and empirical findings will help avoid wasting valuable time and computational budgets during LLM construction.
@@
To stabilize the LLM pre-training, this paper provides theoretical analyses on the spectral norms of the Jacobian matrices for sub-layers to estimate the upper bound of the gradient norm of $\mathcal{L}$. This paper focuses on only the stability of LLM pre-training, and thus, we have to address other issues of LLMs such as hallucinations to use the LLM in a real application.
@@
We do not aim to propose a novel method in this paper, but we mainly focus on theoretical analyses on the spectral norms of the Jacobian matrices to find the factor to stabilize the pre-training of LLMs. We justify our theoretical analyses through experiments with various situations. To activate our modification, we add only several lines to a widely used implementation, i.e., Megatron-LM [^10] https://github.com/NVIDIA/Megatron-LM. Therefore, we believe that it is easy to reproduce our experimental results. Moreover, we did not modify the internal architecture of Transformers. In particular, Scaled Embed only scales embeddings by a constant value, making it convenient in terms of portability across various implementations, such as Transformers in HuggingFace [^11] https://huggingface.co/docs/transformers/main/index and vLLM [^12] https://github.com/vllm-project/vllm.
However, because it is difficult to conclude that our provided conditions completely solve the instability during pre-training of LLMs, it is better to combine other techniques to stabilize the pre-training such as an auxiliary loss described by Chowdhery et al. (2022) to make the pre-training more stable in an actual pre-training situation.
@@ Appendix A Hyper-Parameters
Table 4 shows that hyper-parameters used in our experiments on LLMs. To make the experiments on 13B parameter models close to a realistic situation, we increased the batch size and the number of updates in comparison to ones in experiments on 350M and 1.7B parameter models. In addiiton, for Adam $\beta_{2}$, most studies have used 0.950.950.95 to stabilize their pre-trainings (Brown et al., 2020; Zhang et al., 2022; Zeng et al., 2023; Biderman et al., 2023; Touvron et al., 2023), and thus, we also used 0.950.950.95 in the experiments on 13B parameter models. As described in Section 5.4, we emphasize that the pre-training of Vanilla is essentially unstable even if we use the widely used Adam $\beta_{2}$ value, 0.950.950.95, which is known as the technique to stabilize the pre-training. Therefore, our provided conditions are also effective for the stabilization in the realistic situation.
Appendix B Discussions on Other Configurations
B.1 Methods without Small Sub-Layers
@@
Method Small sub-layers Large shortcut
Xavier Init - -
Vanilla $\checkmark$ -
Embed Detach $\checkmark$ -
Embed LN $\checkmark$ $\checkmark$
Scaled Embed $\checkmark$ $\checkmark$
Xavier Init + Scaled Embed - $\checkmark$
Table 5: Relations between each method in experiments and two conditions to control the upper bound of gradient norms.
@@
Table 6: Perplexities of each method.
B.2 Varying Learning Rate
(a) 350M parameters.
(b) 1.7B parameters.
(a) 350M parameters.
(b) 1.7B parameters.
Le Scao et al. (2022) reported that the stable method, such as Embed LN, was worse than Vanilla. However, in Section 5.3, the stable methods, Scaled Embed and Embed LN, achieved better performance than Vanilla in the 1.7B parameter configuration. We suppose that the difference in the learning rate causes this gap in findings. In this section, we tried to train Vanilla and Scaled Embed with larger and smaller learning rates: $\mathrm{lr}=1.0\times 10^{-3}$ and $1.0\times 10^{-4}$ respectively.
(a) 350M with $\mathrm{lr}=1.0\times 10^{-3}$.
(b) 350M with $\mathrm{lr}=5.0\times 10^{-4}$.
(c) 350M with $\mathrm{lr}=1.0\times 10^{-4}$.
(d) 1.7B with $\mathrm{lr}=1.0\times 10^{-3}$.
(e) 1.7B with $\mathrm{lr}=5.0\times 10^{-4}$.
(f) 1.7B with $\mathrm{lr}=1.0\times 10^{-4}$.
@@ B.2 Varying Learning Rate
[…]
350M parameters
Vanilla 29.96 30.35 34.51 25.12 24.73 32.49
Scaled Embed 28.09 29.86 35.66 22.03 24.37 37.14
1.7B parameters
Vanilla N/A 22.58 23.54 N/A 15.22 16.17
Scaled Embed 20.95 21.29 23.78 12.26 12.53 15.39
B.3 Varying Sequence Length
As described in Section 2.1, Transformers add the output of each sub-layer to the shortcut. Since the standard deviation of the self-attention layer tends to decrease with the length of an input sequence especially at the early stage [^13] See Appendix F for details., a long sequence tends to keep the standard deviation of the shortcut small. Therefore, the long sequence makes the pre-training of Vanilla more unstable.
We conducted experiments with varying the length of the input sequence $L$ from 128128128 to 204820482048. To use the same number of tokens to update parameters, we adjusted the batch size. Figure 9 shows loss values of Vanilla with each $L$ configuration in the validation data. This figure shows that spikes occurred only in the large $L$, i.e., 102410241024 and 204820482048. Moreover, the spikes are more likely to occur at the early stage of the pre-training. Therefore, using a short sequence stabilizes the training at the early stage, as reported in Li et al. (2022).
B.4 RMSNorm
@@
Figure 10 shows the loss values and gradient norms of the Vanilla configuration in Section 5.3 and the one using RMSNorms instead of LNs (“RMSNorm” in figures) with 350M parameters. We trained them with $\mathrm{lr}=5.0\times 10^{-4}$ as in Section 5.3141414We tried to train them with $\mathrm{lr}=1.0\times 10^{-3}$ but RMSNorm exploded.. As shown in these figures, RMSNorm faced loss and gradient spikes in a similar location to the ones of Vanilla. These empirical results also indicate that the RMSNorms have the same problem as LNs regarding the instability.
@@
Appendix C Scaling Embeddings with Larger Value
Equations (15) and (20) indicate that we can stabilize the LLM pre-training by adjusting the standard deviation of the shortcut to a large value. In fact, our experimental results show that we can stabilize the LLM pre-training by making the standard deviation of each embedding close to 111. To investigate how about a larger value, we conducted experiments with making the standard deviation of each embedding close to 5 and 50.
Figure 11 shows loss curves in validation data for 350M and 1.7B parameters in each situation. This figure indicates that although all settings prevented the loss spike problem, the larger standard deviation than 111 degraded the performance. Therefore, it is unnecessary to scale the standard deviation of each embedding with a larger value than 111 to prevent the performance degradation.
@@ Appendix C Scaling Embeddings with Larger Value
Figure 11: Loss curves in validation data when we scale the standard deviation of embeddings with larger than 111.
Appendix D Distributions of Sub-Layer Outputs
Figures 12 and 13 shows output distributions of each sub-layer at the beginning of pre-training and after pre-training. These figures indicate that each sub-layer output is close to a normal distribution in various configurations both at the beginning of pre-training and after pre-training. Therefore, the assumption in this study, which is that the vector $x$ at each layer follows the normal distribution, is reasonable.
@@ Appendix D Distributions of Sub-Layer Outputs
Figure 12: Output distributions of each sub-layer at the beginning of pre-training.
(a) self-attention of 350M parameters.
(b) self-attention of 1.7B parameters.
(c) self-attention of 13B parameters.
(d) FFN of 350M parameters.
(e) FFN of 1.7B parameters.
(f) FFN of 13B parameters.
Figure 13: Output distributions of each sub-layer after pre-training.
Appendix E Discussion on Various Activation Functions in FFN
E.1 ReLU
@@
E.2 SiLU
We consider the case where we use the SiLU function as $\mathcal{F}$. The definition of the SiLU function is as follows:
$\mathrm{SiLU}(x)=x\circ\mathrm{Sigmoid}(x)$ (23)
where $\mathrm{Sigmoid}$ is the sigmoid function.
$\frac{\partial\mathrm{SiLU}(x)}{\partial x}$ $=\mathrm{Sigmoid}(x)+x\circ\frac{\partial\mathrm{Sigmoid}(x)}{\partial x}$
$=\mathrm{Sigmoid}(x)+x\circ\mathrm{Sigmoid}(x)\circ(1-\mathrm{Sigmoid}(x))$
$=\mathrm{Sigmoid}(x)\circ(1+x\circ(1-\mathrm{Sigmoid}(x)))$ (24)
Let $D=\mathrm{diag}\Bigl{(}\frac{\partial\mathrm{SiLU}(W_{1}x)}{\partial W_{1}x}\Bigr{)}\in\mathbb{R}^{d_{\mathrm{ffn}}\times d_{\mathrm{ffn}}}$ Then, we obtain the Jacobian of the $\mathrm{FFN}$ as follows:
$\frac{\partial\mathrm{FFN}(\mathrm{LN}(x^{\prime}))}{\partial\mathrm{LN}(x^{\prime})}=W_{2}DW_{1}$ (25)
Therefore,
$\bigg{\|}\frac{\partial\mathrm{FFN}(\mathrm{LN}(x^{\prime}))}{\partial\mathrm{LN}(x^{\prime})}\bigg{\|}_{2}\leq\|W_{2}\|_{2}\|D\|_{2}\|W_{1}\|_{2}$ (26)
The spectral norm of the diagonal matrix $D$ is the maximum absolute value of its diagonal elements:
$\|D\|_{2}=\max_{i}\bigg{\|}\frac{\partial\text{SiLU}(x_{i})}{\partial x_{i}}\bigg{\|}$ (27)
Moreover, we find that its maximum occurs at $x\approx 2.4$ and is approximately 1.11.11.1, and thus, $\|D\|_{2}\leq 1.1$. Because $\|W_{1}\|_{2}\approx\sigma_{1}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})$ and $\|W_{2}\|_{2}\approx\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})$, we can express the upper bound as the following inequality:
$\bigg{\|}\frac{\partial\mathrm{FFN}(\mathrm{LN}(x^{\prime}))}{\partial\mathrm{LN}(x^{\prime})}\bigg{\|}_{2}\leq 1.1(\sigma_{1}\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})^{2})$ (28)
Finally, Equation (9) can be rewritten as:
$\bigg{\|}\frac{\partial y}{\partial x^{\prime}}\bigg{\|}_{2}$ $\leq 1+1.1\biggl{(}\frac{\sigma_{1}\sigma_{2}}{\sigma_{x^{\prime}}}C_{\mathrm{ffn}}\biggr{)}$ (29)
E.3 SwiGLU
We consider the case where we use the SwiGLU function as $\mathcal{F}$. When we use the SwiGLU function, the FFN layer is expressed as follows:
$\mathrm{FFN}(x)=W_{2}(\mathrm{Swish}(W_{1}x)\circ(Vx))$ (30)
where $V\in\mathbb{R}^{d_{\mathrm{ffn}}\times d}$, and $V$ follows a normal distribution $\mathcal{N}(0,\sigma_{V}^{2})$. Then, we compute the Jacobian of the $\mathrm{FFN}(x)$ as follows:
$\frac{\partial\mathrm{FFN}(x)}{\partial x}=\frac{\partial\mathrm{FFN}(x)}{\partial W_{1}x}\frac{\partial W_{1}x}{\partial x}+\frac{\partial\mathrm{FFN}(x)}{\partial Vx}\frac{\partial Vx}{\partial x}$ (31)
$\frac{\partial\mathrm{FFN}(x)}{\partial W_{1}x}$ $=W_{2}\bigg{(}\mathrm{diag}(Vx)\circ\mathrm{diag}\bigg{(}\frac{\partial\mathrm{Swish}(W_{1}x)}{\partial W_{1}x}\bigg{)}\bigg{)}$ (32)
$\frac{\partial\mathrm{FFN}(x)}{\partial Vx}$ $=W_{2}(\mathrm{diag}(\mathrm{Swish}(W_{1}x)))$ (33)
Therefore, we can rewrite Equation (31) as follows:
$\frac{\partial\mathrm{FFN}(x)}{\partial x}=W_{2}\bigg{(}\mathrm{diag}(Vx)\circ\mathrm{diag}\bigg{(}\frac{\partial\mathrm{Swish(W_{1}x)}}{\partial W_{1}x}\bigg{)}\bigg{)}W_{1}+W_{2}(\mathrm{diag}(\mathrm{Swish}(W_{1}x)))V$ (34)
Let $J_{1}=W_{2}\bigg{(}\mathrm{diag}(Vx)\circ\mathrm{diag}\bigg{(}\frac{\partial\mathrm{Swish}(W_{1}x)}{\partial W_{1}x}\bigg{)}\bigg{)}W_{1}$ and $J_{2}=W_{2}(\mathrm{diag}(\mathrm{Swish}(W_{1}x)))V$, $\frac{\partial\mathrm{FFN}(x)}{\partial x}=J_{1}+J_{2}$. We can derive the upper bound of $\big{\|}\frac{\partial\mathrm{FFN}(x)}{\partial x}\big{\|}_{2}$ as follows:
$\bigg{\|}\frac{\partial\mathrm{FFN}(x)}{\partial x}\bigg{\|}_{2}\leq\|J_{1}\|_{2}+\|J_{2}\|_{2}$ (35)
For $\|J_{1}\|_{2}$, we have:
$\|J_{1}\|_{2}\leq\|W_{2}\|_{2}\|\mathrm{diag}(Vx)\|_{2}\bigg{\|}\mathrm{diag}\bigg{(}\frac{\partial\mathrm{Swish}(W_{1}x)}{\partial W_{1}x}\bigg{)}\bigg{\|}_{2}\|W_{1}\|_{2}$ (36)
Each element of $Vx$ is a sum of $d$ independent random variables with variance $\sigma_{x}^{2}\sigma_{V}^{2}$, and thus, $\mathrm{var}(Vx)=d\sigma_{x}^{2}\sigma_{V}^{2}$. Therefore, from the expected maximum of $d_{\mathrm{ffn}}$ Gaussian random variables,
$\|\mathrm{diag}(Vx)\|_{2}\leq\sigma_{x}\sigma_{V}\sqrt{2\ d\ \mathrm{log}\ d_{\mathrm{ffn}}}$ (37)
The derivation of the $\mathrm{Swish}$ function is bounded by 1.11.11.1 in the same manner as the SiLU function:
$\bigg{\|}\frac{\partial\mathrm{Swish}(W_{1}x)}{\partial W_{1}x}\bigg{\|}_{2}\leq 1.1$ (38)
The spectral norms of $W_{1}$ and $W_{2}$ can be obtained $\|W_{1}\|_{2}\approx\sigma_{1}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})$ and $\|W_{2}\|_{2}\approx\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})$ as described in Section 3.1. We can obtain the upper bound of $\|J_{1}\|_{2}$ with these equations:
$\|J_{1}\|_{2}\leq 1.1\sigma_{x}\sigma_{V}\sigma_{1}\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})^{2}\sqrt{2\ d\ \mathrm{log}\ d_{\mathrm{ffn}}}$ (39)
For $\|J_{2}\|_{2}$, we have:
$\|J_{2}\|_{2}\leq\|W_{2}\|_{2}\|(\mathrm{diag}(\mathrm{Swish}(W_{1}x)))\|_{2}\|V\|_{2}$ (40)
Due to $|\mathrm{Swish}(W_{1}x)|\leq|W_{1}x|$ and $\mathrm{var}(W_{1}x)=d\sigma_{x}^{2}\sigma_{1}^{2}$, we can obtain:
$\|(\mathrm{diag}(\mathrm{Swish}(W_{1}x)))\|_{2}\leq\sigma_{x}\sigma_{1}\sqrt{2\ d\ \mathrm{log}\ d_{\mathrm{ffn}}}$ (41)
Therefore,
$\|J_{2}\|_{2}\leq\sigma_{x}\sigma_{V}\sigma_{1}\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})^{2}\sqrt{2\ d\ \mathrm{log}\ d_{\mathrm{ffn}}}$ (42)
Based on these equations, we can derive the upper bound as follows:
$\bigg{\|}\frac{\partial\mathrm{FFN}(x)}{\partial x}\bigg{\|}_{2}$ $\leq\|J_{1}\|_{2}+\|J_{2}\|_{2}$
$=2.1\ \sigma_{x}\sigma_{V}\sigma_{1}\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})^{2}\sqrt{2\ d\ \mathrm{log}\ d_{\mathrm{ffn}}}$
$=\sigma_{x}\sigma_{V}\sigma_{1}\sigma_{2}C_{\mathrm{swiglu}}$ (43)
where $C_{\mathrm{swiglu}}$ includes $2.1\ (\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})^{2}\sqrt{2\ d\ \mathrm{log}\ d_{\mathrm{ffn}}}$ for the simplification.
Finally, we consider the actual Transformer layer that includes layer normalization and residual connection:
$\bigg{\|}\frac{\partial y}{\partial x^{\prime}}\bigg{\|}_{2}$ $\leq 1+\frac{\sigma_{V}\sigma_{1}\sigma_{2}}{\sigma_{x^{\prime}}}C_{\mathrm{swiglu}}$ (44)
We note that $\sigma_{x}$ in Equation (43) is equal to 111 in the actual Transformer layer because we apply the layer normalization to the input of the FFN layer.